40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4710     Entry time: Fri May 13 01:42:35 2011
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: ready for Schnupp asymmetry measurement 

[Valera / Kiwamu]

 We are able to lock each arm smoothly.  It is ready for the Schnupp asymmetry measurement.

( to be done )

 - Manual D-phase adjustment of the AS55 channel.

 - A script to adjust the D-phase.

 - Trigger logic for the boost filters.

 - Modification of some old alignment scripts to adopt them to the new LSC channels

ELOG V3.1.3-